MAGNETIC TRAPPING OF ^{3}Σ MOLECULES AND TIME-DOMAIN MEASUREMENT OF VIBRATIONAL LIFETIMES

\underline{WESLEY C. CAMPBELL}, MATTHEW T. HUMMON, HSIN-I LU, LAURENS D. VAN BUUREN, EDEM TSIKATA, and JOHN M. DOYLE, Harvard-MIT Center for Ultracold Atoms, Cambridge, MA 02138.

NH molecular radicals are magnetically trapped using helium buffer-gas loading at ∼600 mK. The helium collision Zeeman relaxation and energy transport cross sections are measured, and the ratio of these cross sections is σ_{d}/σ_{in} = 7 × 10^{4}, sufficient for trap lifetimes of order 1 s. The scaling of the helium-induced Zeeman relaxation of^{3}Σ molecules is investigated by changing the radical rotational constant through isotopic substitution. We also measure the spontaneous emission lifetime of NH(X, v = 1), providing a precision measurement of this transition strength for the astrophysical community.